OpenStudy (anonymous):

show that 4n(n^2-1) is divisible by 24 show that 4n(n^2-1) is divisible by 24 @Mathematics

OpenStudy (anonymous):

my approach is as follows: n(n^2-1) is divisible by 6 clearly true for n=2 as that is 2(4-1) = 6 now assume true for n, and try and prove for n+1 expand to n(n-1)(n+1) for n and for n+1 its n(n+1)(n+2) which is n(n+1)(n-1 +3) which is n(n+1)(n-1) + 3n(n+1) now lhs is divisible by 6 as it is the case for n and rhs is an even multiple of 3 which is always divisble by 6, its even because n (n+1) must be an even number now is there a better way than that?

OpenStudy (amistre64):

induction works

OpenStudy (anonymous):

except for the fact that 6 is not divisible by 24!

OpenStudy (anonymous):

should say \[n\geq 3\]

OpenStudy (amistre64):

simplify and work the smaller numbers

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

i'm saying that 4 x z/24 is like saying z/6

OpenStudy (anonymous):

funny the question didn't say n>= 3

OpenStudy (amistre64):

4n(n^2-1); n = 0,1,2,3,... {0,0,24,96, ...}

OpenStudy (anonymous):

yeah i know, i just meant that the whole thing should start with \[n\geq 3\] that is all

OpenStudy (anonymous):

is there a better way to show it than mine

OpenStudy (amistre64):

is n across the reals or the integers?

OpenStudy (anonymous):

integers only

OpenStudy (anonymous):

you have the right idea. you want simply that \[(n-1)n(n+1)\] is divisible by 6, but i doubt you need induction.

OpenStudy (anonymous):

you have three consecutive integers. one is even and one is divisible by 3.

OpenStudy (anonymous):

good point

OpenStudy (anonymous):

thanks all for your help